# Decentralized Uncertainty-Aware Multi-Agent Collision Avoidance With Model Predictive Path Integral

### v1

@@ Abstract
Decentralized multi-agent navigation under uncertainty is a complex task that arises in numerous robotic applications. It requires collision avoidance strategies that account for both kinematic constraints, sensing and action execution noise. In this paper, we propose a novel approach that integrates the Model Predictive Path Integral (MPPI) with a probabilistic adaptation of Optimal Reciprocal Collision Avoidance. Our method ensures safe and efficient multi-agent navigation by incorporating probabilistic safety constraints directly into the MPPI sampling process via a Second-Order Cone Programming formulation. This approach enables agents to operate independently using local noisy observations while maintaining safety guarantees. We validate our algorithm through extensive simulations with differential-drive robots and benchmark it against state-of-the-art methods, including ORCA-DD and B-UAVC. Results demonstrate that our approach outperforms them while achieving high success rates, even in densely populated environments. Additionally, validation in the Gazebo simulator confirms its practical applicability to robotic platforms.
@@ IV-A Model Predictive Path Integral
Let $\mathcal{V}=(\mathbf{\nu_{0}},\mathbf{\nu_{1}},...,\mathbf{\nu_{T-1}})$ be the control sequence, $\mathcal{U}$ the set of admissible controls, and $X=(\mathbf{x_{0}},\mathbf{x_{1}},…,\mathbf{x_{T}})$ the state trajectory over horizon $T$. The cost function is:
@@
Given the current state $\mathbf{x_{0}}$ and an initial control sequence $U^{init}=(\mathbf{u^{init}_{0}},…,\mathbf{u^{init}_{T-1}})$. The MPPI algorithm samples $K$ noise sequences $\xi^{k}=(\mathbf{\epsilon^{k}_{0}},…,\mathbf{\epsilon^{k}_{T-1}}),\;\mathbf{\epsilon^{k}_{t}}\sim\mathcal{N}(0,\Sigma^{*})$. Notably that the sampling variance $\Sigma^{*}$ may be higher than the natural variance $\Sigma$ (but also should be diagonal). Based on $U^{init}$ and $\xi^{k}$, a set of $K$ control sequences $U^{k}$ is obtained:
$U^{k}=(\mathbf{u^{k}_{0}},…,\mathbf{u^{k}_{T-1}}),\quad\mathbf{u^{k}_{t}}=\mathbf{u^{init}_{t}}+\mathbf{\epsilon^{k}_{t}}$ (6)
@@ IV-B Optimal Reciprocal Collision Avoidance
$VO^{\tau}_{i|j}=\{\mathbf{v}\;|\;\exists t’\in[0,\tau]:t^{\prime}\cdot\mathbf{v}\in D(\mathbf{p_{t}^{j}}-\mathbf{p_{t}^{i}},r^{i}+r^{j})\}$ (11)
@@ VI EXPERIMENTAL EVALUATION
The proposed method was implemented in C++ [^1] http://github.com/PathPlanning/MPPI-Collision-Avoidance and evaluated in various experimental setups. Initially, experiments were conducted using a widely adopted model of differential-drive robot motion. This system model has been extensively studied in the literature on collision avoidance with kinematic constraints [3, 11, 6], enabling a comparative analysis with several well-established methods. Subsequently, the algorithm was integrated into the ROS2 Navigation stack as a controller module and validated using a well-known Gazebo robotic simulator.
@@ VI-A1 Experimental Setup
Figure 3: The average makespan values and its standard deviation for evaluated algorithms for different number of agents in Circle and Random scenarios. Each point on the graph is included only if more than 50% of the runs were successfully completed. The lower is the better.
@@ VI-A2 Experimental Results
The resultant success rates are reported in Table I as well as percentage of runs that were terminated due to agent collisions. Evidently, the proposed approach successfully handled all tasks in Circle scenario without any collisions. In Random scenario the proposed algorithm solved the highest percentage of tasks, allowing collisions in only one run, where the occurrence of collisions can be explained by the fact that our method provides safety guarantees in the probabilistic sense and the imposed safety threshold is close to 1 but not 1. In contrast, its predecessor, MPPI-ORCA is highly prone to collisions. ORCA-DD and B-UAVC exhibit lower collision rates but often result in deadlocks when the robots get safely stuck and do not progress towards the goal.
@@ VI-B Simulation in Gazebo
Additionally, we evaluated the Dynamic Window Approach (DWA) [36], which is also implemented within the ROS2 Navigation Stack. However, this algorithm also failed to successfully complete the task. For further details, please refer to the publicly-available video [^2] https://youtu.be/_D4zDYJ4KCk.

### v2

@@ Abstract
Decentralized multi-agent navigation under uncertainty is a complex task that arises in numerous robotic applications. It requires collision avoidance strategies that account for both kinematic constraints, sensing and action execution noise. In this paper, we propose a novel approach that integrates the Model Predictive Path Integral (MPPI) with a probabilistic adaptation of Optimal Reciprocal Collision Avoidance. Our method ensures safe and efficient multi-agent navigation by incorporating probabilistic safety constraints directly into the MPPI sampling process via a Second-Order Cone Programming formulation. This approach enables agents to operate independently using local noisy observations while maintaining safety guarantees. We validate our algorithm through extensive simulations with differential-drive robots and benchmark it against state-of-the-art methods, including ORCA-DD and B-UAVC. Results demonstrate that our approach outperforms them while achieving high success rates, even in densely populated environments. Additionally, validation in the Gazebo simulator confirms its practical applicability to robotic platforms. A source code is available at: http://github.com/PathPlanning/MPPI-Collision-Avoidance.
@@ IV-A Model Predictive Path Integral
Let $\mathcal{V}=(\mathbf{\nu_{0}},\mathbf{\nu_{1}},\ldots,\mathbf{\nu_{T-1}})$ be the control sequence, $\mathcal{U}$ the set of admissible controls, and $X=(\mathbf{x_{0}},\mathbf{x_{1}},\ldots,\mathbf{x_{T}})$ the state trajectory over horizon $T$. The cost function is:
@@
Given the current state $\mathbf{x_{0}}$ and an initial control sequence $U^{init}=(\mathbf{u^{init}_{0}},\ldots,\mathbf{u^{init}_{T-1}})$. The MPPI algorithm samples $K$ noise sequences $\xi^{k}=(\mathbf{\epsilon^{k}_{0}},\ldots,\mathbf{\epsilon^{k}_{T-1}}),\;\mathbf{\epsilon^{k}_{t}}\sim\mathcal{N}(0,\Sigma^{*})$. Notably that the sampling variance $\Sigma^{*}$ may be higher than the natural variance $\Sigma$ (but also should be diagonal). Based on $U^{init}$ and $\xi^{k}$, a set of $K$ control sequences $U^{k}$ is obtained:
$U^{k}=(\mathbf{u^{k}_{0}},\ldots,\mathbf{u^{k}_{T-1}}),\quad\mathbf{u^{k}_{t}}=\mathbf{u^{init}_{t}}+\mathbf{\epsilon^{k}_{t}}$ (6)
@@ IV-B Optimal Reciprocal Collision Avoidance
$VO^{\tau}_{i|j}=\{\mathbf{v}\;|\;\exists t^{\prime}\in[0,\tau]:t^{\prime}\cdot\mathbf{v}\in D(\mathbf{p_{t}^{j}}-\mathbf{p_{t}^{i}},r^{i}+r^{j})\}$ (11)
@@ VI EXPERIMENTAL EVALUATION
The proposed method was implemented in C++ and evaluated in various experimental setups. Initially, experiments were conducted using a widely adopted model of differential-drive robot motion. This system model has been extensively studied in the literature on collision avoidance with kinematic constraints [3, 11, 6], enabling a comparative analysis with several well-established methods. Subsequently, the algorithm was integrated into the ROS2 Navigation stack as a controller module and validated using a well-known Gazebo robotic simulator.
@@ VI-A1 Experimental Setup
Figure 3: The average makespan for evaluated algorithms for different number of agents in Circle and Random scenarios. Each point on the graph is included only if more than 50% of the runs were successful. The lower is the better.
@@ VI-A2 Experimental Results
The resultant success rates are reported in Table I as well as the percentage of runs that were terminated due to agent collisions. Evidently, the proposed approach successfully handled all tasks in both the Circle and Random scenarios, demonstrating strong safety guarantees. In contrast, its predecessor, MPPI-ORCA is highly prone to collisions. ORCA-DD and B-UAVC exhibit lower collision rates but often result in deadlocks when the robots get safely stuck and do not progress towards the goal.
@@ VI-B Simulation in Gazebo
Additionally, we evaluated the Dynamic Window Approach (DWA) [36], which is also implemented within the ROS2 Navigation Stack. However, this algorithm also failed to successfully complete the task. For further details, please refer to the publicly-available supplementary video [^1] https://youtu.be/_D4zDYJ4KCk.